# Degrees of the Wasserstein Distance to Small Toric Models

### v1

@@ 1. Introduction
$\Delta_{n-1}\,:=\,\{(p_{1},\ldots,p_{n})\,:\,\sum_{i=1}^{n}p_{i}=1,\mbox{ and }p_{i}\geq 0,\,i=1,\ldots,n\}$
consists of probability distributions of a discrete random variable with a state space of size $n$. We take this state space to be $[n]:=\{1,\ldots,n\}$. A statistical model $\mathcal{M}$ is a subset of $\Delta_{n-1}$ which represents distributions to which a hypothesized unknown distribution $\nu$ belongs. Typically, after collecting data $u=(u_{1},\ldots,u_{n})$ where $u_{i}$ is the number of times outcome $i$ is observed, one forms the empirical distribution ${\overline{\mu}}=\frac{1}{N}u$ where $N=\sum_{i=1}^{n}u_{i}$ is the sample size. Note that ${\overline{\mu}}\in\Delta_{n-1}$. To estimate the unknown distribution $\nu$ a standard approach is to locate $\nu\in\mathcal{M}$ that is a “closest” point to ${\overline{\mu}}$. For instance, $\nu$ can be taken to be the maximum likelihood estimator [17, Chapter 7] of ${\overline{\mu}}$. In this case, $\nu$ is the point on $\mathcal{M}$ that minimizes the Kullback-Leibler divergence from ${\overline{\mu}}$ to $\mathcal{M}$. However, Kullback-Leibler divergence is not a metric, and the maximum likelihood estimator does not minimize a true distance function from ${\overline{\mu}}$ to $\mathcal{M}$.
For the above density estimation problem, one can use a distance minimization approach if the state space $[n]$ is also a metric space. A metric on $[n]$ is a collection of nonnegative real numbers $d_{ij}$ for $i,j\in[n]$ such that $d_{ii}=0$ for all $i\in[n]$, $d_{ij}=d_{ji}$, and the triangle inequality $d_{ik}\leq d_{ij}+d_{jk}$ holds for all $i,j,k\in[n]$.
@@
that contains $(1,1,\ldots,1)$ in its row space where $\bf a_{j}$ denotes the $jth$ column of $A$, $j=1,2,\ldots,n$. Then $X$ is the Zariski closure of the image of the monomial map
(*$(\mathbb{C}^{*})^{d}$ (*$\longrightarrow(\mathbb{C}^{*})^{n}$
@@
We use the discrete metric on $[4]=\{1,2,3,4\}$. With this, the Wasserstein ball is a cuboctahedron with $12$ vertices, $24$ edges, and $14$ facets. Figure 1 is a picture of the simplex $\Delta_{3}$ together with the twisted cubic (orange) curve and the Wasserstein ball centered at $\mu=\left(\frac{1}{4},\frac{1}{5},\frac{1}{6},\frac{23}{60}\right)$. Slowly the radius $\lambda$ of the Wasserstein ball around $\mu$ is decreased until it touches the curve. The snapshot shows $\lambda\approx 0.268$ when the ball just touches the curve at the red point $\nu\approx(0.02,0.16,0.44,0.38)$, and the point lies on an edge of the Wasserstein ball.
@@
Understanding the geometry of the distance minimization problem from a point to an algebraic variety is part of metric algebraic geometry [2, 19]. This problem has been studied extensively when the distance metric is the usual Euclidean metric leading to the Euclidean distance degree of an algebraic variety; see [7], [11]. We wish to continue studying this problem in the case of the Wasserstein metric first considered in [4].
@@
[…]
@@ 2.1. Polar Degrees
Let $i\in\{0,1,\ldots,\dim(X)\}$. If $V$ is generic with $\dim V=\text{codim}X-2+i$, then the degree, $\deg(P(X,V))$, of the polar variety $P(X,V)$ is independent of $V$.
@@
The integer $\mu_{i}(X)$, for $i=\dim V-\text{codim }X+2$, is called the $i^{th}$ polar degree of $X$, where
@@
[…]
Notice that Proposition 2.3 can be used to assign a natural bound on what subspaces $V$ we need to consider when computing polar degrees of $X$. When $\dim(V)$ is small enough, the non-transversality condition will hold for all points of $X$, hence $P(X,V)=X$. Therefore, we use those $V$ which satisfy $\dim(V)=\text{codim}(X)-2+i$. In other words, we consider $V$ where $\text{codim}(X)-2\leq\dim(V)\leq n-2$.
@@
$\text{codim}(X)-2\leq\dim(V)\leq n-2\implies 1\leq\dim(V)\leq 2.$
  • If $V$ has dimension $1$, from the relation $\dim(V)=\text{codim}(X)-2+i,$ we have $i=0$, and by [2, Theorem 4.14(b)]
@@
It is known that $\delta_{j}(X)=0$ whenever $j<\text{codim}(X^{\vee})$ or $j>\dim(X)+1$. Moreover, for $j=\text{codim}(X^{\vee})$ and $j=\dim(X)+1$, $\delta_{j}(X)=\deg(X^{\vee})$ and $\delta_{j}(X)=\deg(X)$, respectively.
@@ 2.2. Rational Normal Scrolls
A rational normal scroll is a toric variety associated to a $(d+1)\times n$ matrix $A$ of the following form:
@@
The rational normal scroll $S=S(n_{1},\ldots,n_{d})$ is equal to the intersection of the Segre embedding $X_{1,N-1}$ of $\mathbb{P}^{1}\times\mathbb{P}^{N-1}$ in $\mathbb{P}^{2N-1}$ with a linear subspace $\Pi$ of dimension $(n-1)$. Therefore $S^{\vee}$ is obtained by projecting $X_{1,N-1}^{\vee}$ from $\Pi^{\vee}$. It is known that $X_{1,N-1}^{\vee}=X_{1,N-1}$. Since the degree of $X_{1,N-1}$ is $N$ we conclude that $\deg(S^{\vee})=N$. This also means that $\text{codim}(S^{\vee})=(n-1)-N=d-1$. Therefore $\delta_{d-1}=N$. Because $\deg(S)$ is also $N$, for $\dim(S)+1=d+1$ we get $\delta_{d+1}=N$. Furthermore, according to [2, Theorem 4.15] $\mu_{1}(S^{\vee})=\mu_{1}(S)=\delta_{d}(S)$ which means that $\mu_{1}(S^{\vee})=\delta_{N}(S^{\vee})$. Again, since $S^{\vee}$ is obtained by projecting $X_{1,N-1}^{\vee}=X_{1,N-1}$ from $\Pi^{\vee}$, we need to compute $\delta_{N}(X_{1,N-1})$. We use the formula from [4, Corollary 16] to compute this polar degree to be $2(N-1)$. ∎
@@ 2.3. Small graphical models
where $I\in\prod_{i}^{n}[d_{i}]$. When $I=(i_{1},\ldots,i_{n})$ and $V(C)=\{j_{1},\ldots,j_{k}\}$ we let $I(C)=(i_{j_{1}},i_{j_{2}},\ldots,i_{j_{k}})$. The toric variety $X_{G,\mathbf{d}}$ is a projective variety in $\mathbb{P}^{D-1}$ where $D=\prod_{i=1}^{n}d_{i}$. The graphical model $\mathcal{M}_{G,\mathbf{d}}$ is equal to $X_{G,\mathbf{d}}\cap\Delta_{D-1}$. If $G$ consists of $n$ isolated vertices $\mathcal{M}_{G,\mathbf{d}}$ is called a complete independence model. In this case, $X_{G,\mathbf{d}}$ is the variety of all $d_{1}\times d_{2}\times\cdots\times d_{n}$–tensors of rank one.
@@ 2.3.1. Polar degrees of star trees
We label the leaves of the tree by $1,\ldots,n$ and let the label of the internal vertex be $n+1$. The resulting model $\mathcal{M}_{G,\mathbf{d}}$ is a conditional independence model corresponding to the conditional independence statement $Z_{1}\,\perp\,Z_{2}\,\perp\cdots\,\perp\,Z_{n}\mid Z_{n+1}$ for all $1\leq i<j\leq n$. The equations defining $X_{G,\mathbf{d}}$ in this case are easy to describe:
@@
Let $G$ be the star tree with two leaves where the two random variables corresponding to them are binary. We consider the cases where the random variable for the internal node is binary and ternary. Here, $Y_{2,(2,2)}$ is the Segre embedding of $\mathbb{P}\times\mathbb{P}$. The multidegree of the conormal variety of $Y_{2,(2,2)}$ is
@@ 2.3.3. Polar degrees of the binary $4$-cycle
In all the examples we have looked at so far, we dealt with toric varieties associated to graphical models that are decomposable [17] [12]. Our next example is the smallest graphical model that is not decomposable. The underlying graph $G$ is a cycle with four vertices. We take all four random variables to be binary. The toric variety $X_{G,(2,2,2,2)}$ has the parametrization
@@ 2.3.4. Polar degrees of the no-three-way interaction models
The no-three-way interaction model is a toric model based on the graph $G$ that is a triangle, though it is not a graphical model. It belongs to a larger class of models known as hierarchical log-linear models; see [17]. If the three random variables have state spaces $[r]$, $[s]$, and $[t]$, the model is parametrized by
@@
{tikzpicture}
Figure 6. The no-three-way interaction model.
We were able to compute the polar degrees when $r=s=t=2$ and $r=s=2,\,t=3$. The first variety is a hypersurface in $\mathbb{P}^{7}$ of degree $4$. Its polar degrees can be read off from
@@ 3. Computing Wasserstein Degrees
algebraically. Here $\mathcal{M}=X\cap\Delta_{n-1}$ is a toric model, and $F$ is a face of the Wasserstein ball $B$ for a Wasserstein distance based on a metric on $[n]$. The objective function $\ell_{F}$ is a linear functional that attains its maximum on $F$ over $B$. A natural choice is given by a normal vector $n$ to the face $F$ and defining $\ell_{F}(\nu):=n^{\top}\nu$. Finally, $\mu+\mathcal{L}_{F}$ is the affine linear subspace that is a translate of $\mathcal{L}_{F}$ by a generic point $\mu\in\Delta_{n-1}$, where $\mathcal{L}_{F}$ is the subspace spanned by the vertices of $F$. This is a constrained optimization problem with two types of constraints: linear constraints defining the affine space $\mu+\mathcal{L}_{F}$ and polynomials defining $\mathcal{M}$; here, we are ignoring the non-negativity constraints on the coordinate variables. Since the feasible set is non-convex, we solve this problem by computing the complex critical points of $\ell_{F}$ on $(\mu+\mathcal{L}_{F})\cap\mathcal{M}$. Recall that we defined the Wasserstein degree $w(X,F)$ of $X$ with respect to $F$ as the number of these complex critical points for generic $\mu\in\Delta_{n-1}$.
One way of computing these complex critical points is by introducing Lagrange multipliers. This adds extra variables, as many as there are constraints. Instead, we use an equivalent formulation without the extra Lagrange multipliers, however, at the expense of introducing more equations. We now explain this.
@@
$c=\text{codim}(I)+1$. Then, the first condition above holds if and only if the augmented Jacobian matrix evaluated at $\nu^{*}$,
@@ 3.1. Computational experiments
The augmented Jacobian matrix $\overline{\mathrm{Jac}}(p)$ is a matrix where almost all the rows are given by the polynomials $g_{1},\ldots,g_{m}$ defining $\mathcal{M}$. These can grow quickly, and the number of $c$-minors where $c=\text{codim}(\mathcal{M})+1$ is on the order of $\binom{n}{c}\binom{m}{c}$. The computations need to be repeated for each face of $B$. As a result, the models we could consider were limited.
@@
In the following tables, we report the results of our computations for the binary path model on $3$ vertices, the binary no-three-way interaction model, and various Hirzebruch surfaces. For the first two, the distance metric is the Hamming distance on $[2]\times[2]\times[2]$. For the last, it is the $L_{1}$-metric on $[n]$ where $n=a+b+2$ for the Hirzebruch surface $S(a,b)$. We report the frequency of the Wasserstein degrees observed for each model and each $i$-dimensional face of the corresponding Wasserstein ball $B$. Each table contains the $f$-vector $(f_{0},f_{1},\ldots)$ of $B$ where $f_{i}$ is the number of $i$-dimensional faces.

### v2

@@ 1. Introduction
$\Delta_{n-1}\,:=\,\left\{(p_{1},\ldots,p_{n})\,:\,\sum_{i=1}^{n}p_{i}=1,\mbox{and }p_{i}\geq 0,\,i=1,\ldots,n\right\}$
consists of probability distributions of a discrete random variable with a state space of size $n$. We take this state space to be $[n]:=\{1,\ldots,n\}$. A statistical model $\mathcal{M}$ is a subset of $\Delta_{n-1}$ which represents distributions to which a hypothesized unknown distribution $\nu$ belongs. Typically, after collecting data $u=(u_{1},\ldots,u_{n})$ where $u_{i}$ is the number of times outcome $i$ is observed, one forms the empirical distribution ${\overline{\mu}}=\frac{1}{N}u$ where $N=\sum_{i=1}^{n}u_{i}$ is the sample size. Note that ${\overline{\mu}}\in\Delta_{n-1}$. To estimate the unknown distribution $\nu$, a standard approach is to locate $\nu\in\mathcal{M}$, that is a “closest” point to ${\overline{\mu}}$. For instance, $\nu$ can be taken to be the maximum likelihood estimator [17, Chapter 7] of ${\overline{\mu}}$. In this case, $\nu$ is the point on $\mathcal{M}$ that minimizes the Kullback-Leibler divergence from ${\overline{\mu}}$ to $\mathcal{M}$. However, Kullback-Leibler divergence is not a metric, and the maximum likelihood estimator does not minimize a true distance function from ${\overline{\mu}}$ to $\mathcal{M}$.
For the above density estimation problem, one can use a distance minimization approach if the state space $[n]$ is also a metric space. A metric on $[n]$ is a collection of nonnegative real numbers $d_{ij}$ for $i,j\in[n]$ such that $d_{ii}=0$ for all $i\in[n]$, $d_{ij}=d_{ji}$, and the triangle inequality $d_{ik}\leq d_{ij}+d_{jk}$ holds for all $i,j,k\in[n]$. Sometimes, the metric on $[n]$ is written as an $n\times n$ nonnegative symmetric matrix $d=(d_{ij})_{i,j\in[n]}$. Common examples include the discrete metric, the $L_{1}$ metric, the $L_{0}$ metric, and the Hamming distance metric.
@@
that contains $(1,1,\ldots,1)$ in its row space where ${\bf a}_{j}$ denotes the $j^{\text{th}}$ column of $A$, $j\in[n]$. Then $X$ is the Zariski closure of the image of the monomial map
$(\mathbb{C}^{*})^{d}$ $\longrightarrow(\mathbb{C}^{*})^{n}$
@@
We use the discrete metric on $[4]=\{1,2,3,4\}$. This is the metric associated with $(d_{12},d_{13},d_{14},d_{23},d_{24},d_{34})=(1,1,\cdots,1)$. With this, the Wasserstein ball is a cuboctahedron with $12$ vertices, $24$ edges, and $14$ facets. Figure 1 is a picture of the simplex $\Delta_{3}$ together with the twisted cubic (orange) curve and the Wasserstein ball (blue) centered at $\mu=\left(\frac{1}{4},\frac{1}{5},\frac{1}{6},\frac{23}{60}\right)$. Slowly the radius $\lambda$ of the Wasserstein ball around $\mu$ is decreased until it touches the curve. The snapshot shows $\lambda\approx 0.268$ when the ball just touches the curve at the red point $\nu\approx(0.02,0.16,0.44,0.38)$, and the point lies on an edge of the Wasserstein ball.
@@
Understanding the geometry of the distance minimization problem from a point to an algebraic variety is part of metric algebraic geometry [2, 19]. This problem has been studied extensively when the distance metric is the usual Euclidean metric leading to the Euclidean distance degree of an algebraic variety; see [7], [11]. We wish to continue studying this problem in the case of the Wasserstein metric, first considered in [4].
@@
$\operatorname{codim}(F)=r+1$ is at most $\mu_{i}(X)$, the $i$th polar degree of $X$ where $i=\dim(X)-r$.
@@ 2.1. Polar Degrees
Let $i\in\{0,1,\ldots,\dim(X)\}$. If $V$ is generic with $\dim(V)=\operatorname{codim}(X)-2+i$, then the degree, $\deg(P(X,V))$, of the polar variety $P(X,V)$ is independent of $V$.
@@
The integer $\mu_{i}(X)$, for $i=\dim(V)-\operatorname{codim}(X)+2$, is called the $i^{th}$ polar degree of $X$, where
@@
$\dim(V)-\operatorname{codim}(X)<\dim(T_{p}X\cap V).$ (7)
Notice that Proposition 2.3 can be used to assign a natural bound on what subspaces $V$ we need to consider when computing polar degrees of $X$. When $\dim(V)$ is small enough, the non-transversality condition will hold for all points of $X$, hence $P(X,V)=X$. Therefore, we use those $V$ which satisfy $\dim(V)=\operatorname{codim}(X)-2+i$. In other words, we consider $V$ where $\operatorname{codim}(X)-2\leq\dim(V)\leq n-2$.
@@
$\operatorname{codim}(X)-2\leq\dim(V)\leq n-2\implies 1\leq\dim(V)\leq 2.$
  • If $V$ has dimension $1$, from the relation $\dim(V)=\operatorname{codim}(X)-2+i,$ we have $i=0$, and by [2, Theorem 4.14(b)]
@@
Recall that $u$ is tangent to $X$ at $x$ if and only if $H_{u}\supset T_{x}X$, where $H_{u}$ is the hyperplane in $\mathbb{P}^{n}$ associated with the point $u\in(\mathbb{P}^{n})^{\vee}$.
@@
It is known that $\delta_{j}(X)=0$ whenever $j<\operatorname{codim}(X^{\vee})$ or $j>\dim(X)+1$. Moreover, for $j=\operatorname{codim}(X^{\vee})$ and $j=\dim(X)+1$, $\delta_{j}(X)=\deg(X^{\vee})$ and $\delta_{j}(X)=\deg(X)$, respectively.
@@ 2.2. Rational Normal Scrolls
A rational normal scroll is a toric variety associated with a $(d+1)\times n$ matrix $A$ of the following form:
@@
The rational normal scroll $S=S(n_{1},\ldots,n_{d})$ is equal to the intersection of the Segre embedding $X_{1,N-1}$ of $\mathbb{P}^{1}\times\mathbb{P}^{N-1}$ in $\mathbb{P}^{2N-1}$ with a linear subspace $\Pi$ of dimension $(n-1)$. Therefore $S^{\vee}$ is obtained by projecting $X_{1,N-1}^{\vee}$ from $\Pi^{\vee}$. It is known that $X_{1,N-1}^{\vee}=X_{1,N-1}$. Since the degree of $X_{1,N-1}$ is $N$ we conclude that $\deg(S^{\vee})=N$. This also means that $\operatorname{codim}(S^{\vee})=(n-1)-N=d-1$. Therefore $\delta_{d-1}=N$. Because $\deg(S)$ is also $N$, for $\dim(S)+1=d+1$ we get $\delta_{d+1}=N$. Furthermore, according to [2, Theorem 4.15] $\mu_{1}(S^{\vee})=\mu_{1}(S)=\delta_{d}(S)$ which means that $\mu_{1}(S^{\vee})=\delta_{N}(S^{\vee})$. Again, since $S^{\vee}$ is obtained by projecting $X_{1,N-1}^{\vee}=X_{1,N-1}$ from $\Pi^{\vee}$, we need to compute $\delta_{N}(X_{1,N-1})$. We use the formula from [4, Corollary 16] to compute this polar degree to be $2(N-1)$. ∎
@@ 2.3. Small graphical models
where $I\in\prod_{i}^{n}[d_{i}]$. When $I=(i_{1},\ldots,i_{n})$ and $V(C)=\{j_{1},\ldots,j_{k}\}$ we let $I(C)=(i_{j_{1}},i_{j_{2}},\ldots,i_{j_{k}})$. The toric variety $X_{G,\mathbf{d}}$ is a projective variety in $\mathbb{P}^{D-1}$ where $D=\prod_{i=1}^{n}d_{i}$. The graphical model $\mathcal{M}_{G,\mathbf{d}}$ is equal to $X_{G,\mathbf{d}}\cap\Delta_{D-1}$. If $G$ consists of $n$ isolated vertices $\mathcal{M}_{G,\mathbf{d}}$ is called a complete independence model.
@@ 2.3.1. Polar degrees of star trees
[…]
@@
Let $G$ be the star tree with two leaves where the two random variables corresponding to them are binary. We consider the cases where the random variable for the internal node is binary and ternary. Here, $Y_{2,(2,2)}$ is the Segre embedding of $\mathbb{P}^{1}\times\mathbb{P}^{1}$. The multidegree of the conormal variety of $Y_{2,(2,2)}$ is
@@ 2.3.3. Polar degrees of the binary $4$-cycle
In all the examples we have looked at so far, we dealt with toric varieties associated with graphical models that are decomposable [17] [12]. Our next example is the smallest graphical model that is not decomposable. The underlying graph $G$ is a cycle with four vertices. We take all four random variables to be binary. The toric variety $X_{G,(2,2,2,2)}$ has the parametrization
@@ 2.3.4. Polar degrees of the no-three-way interaction models
The no-three-way interaction model is a toric model based on the graph $G$, which is a triangle, though it is not a graphical model. It belongs to the larger class of hierarchical log-linear models; see [17]. If the three random variables have state spaces $[r]$, $[s]$, and $[t]$, the model is parametrized by
@@
$X$$Z$$Y$
Figure 6. Graph of the no-three-way interaction model.
We were able to compute the polar degrees when $r=s=t=2$ and $r=s=2,\,t=3$. The first variety is a hypersurface in $\mathbb{P}^{7}$ of degree $4$. Its polar degrees can be read from
@@ 3. Computing Wasserstein Degrees
algebraically. Here $\mathcal{M}=X\cap\Delta_{n-1}$ is a toric model, and $F$ is a face of the Wasserstein ball $B$ for a Wasserstein distance based on a metric on $[n]$. The objective function $\ell_{F}$ is a linear functional that attains its maximum on $F$ over $B$. A natural choice is given by a normal vector $n$ to the face $F$ and defining $\ell_{F}(\nu):=n^{\top}\nu$. Finally, $\mu+\mathcal{L}_{F}$ is the affine linear subspace that is a translation of $\mathcal{L}_{F}$ by a generic point $\mu\in\Delta_{n-1}$, where $\mathcal{L}_{F}$ is the subspace spanned by the vertices of $F$. This is a constrained optimization problem with two types of constraints: linear constraints defining the affine space $\mu+\mathcal{L}_{F}$ and polynomials defining $\mathcal{M}$; here, we are ignoring the non-negativity constraints on the coordinate variables. Since the feasible set is non-convex, we solve this problem by computing the complex critical points of $\ell_{F}$ on $(\mu+\mathcal{L}_{F})\cap\mathcal{M}$. Recall that we defined the Wasserstein degree $w(X,F)$ of $X$ with respect to $F$ as the number of these complex critical points for generic $\mu\in\Delta_{n-1}$.
One way of computing these complex critical points is by introducing Lagrange multipliers. This adds extra variables, as many as there are constraints. Instead, we use an equivalent formulation without the extra Lagrange multipliers, but at the expense of introducing more equations. We now explain this.
@@
$c=\operatorname{codim}(I)+1$. Then, the first condition above holds if and only if the augmented Jacobian matrix evaluated at $\nu^{*}$,
@@ 3.1. Computational experiments
The augmented Jacobian matrix $\overline{\mathrm{Jac}}(p)$ is a matrix where almost all the rows are given by the polynomials $g_{1},\ldots,g_{m}$ defining $\mathcal{M}$. These can grow quickly, and the number of $c$-minors where $c=\operatorname{codim}(\mathcal{M})+1$ is on the order of $\binom{n}{c}\binom{m}{c}$. The computations need to be repeated for each face of $B$. As a result, the models we could consider were limited.
@@
In the following tables, we report the results of our computations for the binary path model on $3$ vertices, the binary no-three-way interaction model, and various Hirzebruch surfaces. For the first two, the distance metric is the Hamming distance on $[2]\times[2]\times[2]$. That is,
[…]
@@
We would like to point out that, ultimately, the real critical points are most relevant to solve the Wasserstein distance problem. The number of the real critical points will stay constant for points $\mu$ in certain regions of the probability simplex $\Delta_{n-1}$. These regions should be the connected regions of the complement of a discriminant hypersurface, reminiscent of the Euclidean distance degree discriminants [7]. We leave this to a future project.
Acknowledgements
We thank Bernd Sturmfels and Jose Israel Rodriguez for their helpful conversations. We also appreciate the valuable comments of the anonymous referee.